### v1

# Guided simulation of conditioned chemical reaction networks

Abstract.
Let $X$ be a chemical reaction process, modeled as a multi-dimensional continuous-time jump process. Assume that at given times $0<t_{1}<\cdots<t_{n}$, linear combinations $v_{i}=L_{i}X(t_{i}),\,i=1,\dots,n$ are observed for given matrices $L_{i}$. We show how to sample the process conditioned on hitting the states $v_{1},\dots,v_{n}$ by a change of measure on the law of the unconditioned process. We derive sufficient conditions for this change of measure and complement our results using numerical illustrations.   Keywords: Chemical reaction processes, Doob’s $h$-transform, exponential change of measure, guided process.   AMS subject classification: 60J27, 60J28, 60J74
footnotetext: Contact: M.A.Corstanje@vu.nl
1. Introduction
Chemical reaction networks are used to study a wide class of biological, physical and chemical processes that evolve over time. For instance, one can think of the transcription of genes to mRNA and then the translation to protein, the kinetics of a virus or the dynamics of chemical components reacting with each other. The forward evolution of such processes can be described in different ways: (i) a system of ordinary differential equations, see e.g. Feinberg, (2019), (ii) a system of stochastic differential equations, see e.g. Fuchs, (2013) or (iii) continuous-time Markov jump processes, as in Anderson, (2007). It is the latter option that we consider in this paper.
1.1. Chemical reaction networks
Chemical reactions are described as linear combinations of chemical components merging into each other. Typically, one denotes a reaction in which components $A$ and $B$ are merged into $C$ and $D$ by
$A+B\to C+D$
More generally, a chemical reaction network consists of:
  • a Species set ${\mathcal{S}}=\left\{S_{1},\dots,S_{d}\right\}$ which consists of the chemical components whose counts we model;
  • a Reaction set ${\mathcal{R}}$. A reaction $\ell\in{\mathcal{R}}$ is modeled as $\sum_{k}\nu_{k\ell}S_{k}\to\sum_{k}\nu_{k\ell}^{\prime}S_{k}$. We characterize $\ell$ by the change of counts of the species $\xi_{\ell}=\left(\nu_{k\ell}^{\prime}-\nu_{k\ell}\right)_{k=1}^{n}$.
Let ${\mathbb{S}}\subseteq{\mathbb{Z}}_{+}^{d}$ denote the state space of the reaction system. A vector in ${\mathbb{S}}$ is of the form $(x_{k})_{k=1}^{d}$, where $x_{k}$ denotes the species count of type $S_{k}$. We study a Markov process $X=\left(X(t)\right)_{t\geq 0}$ on ${\mathbb{S}}$ that models the evolution of species counts over time. That is, at time $t$, the $k$-th element of $X(t)$ is given by #$\#S_{k}(t)$. At a given time $t$, the time until a reaction of type $\ell$ takes place is given by $\tau_{\ell}$. When reaction ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ occurs, the process jumps at time $t+\tau_{\hat{\ell}}$ to $X\left(t+\tau_{\hat{\ell}}\right)=X(t)+\xi_{\hat{\ell}}$. A more detailed description of the stochastic model is given in Section 2. The process $X$ evolving on the chemical reaction network is referred to as the chemical reaction process.
1.2. Statistical problem
Suppose at fixed times $0=t_{0}<t_{1}<\cdots<t_{n}$, we observe $v_{1},\dots v_{n}$, where $v_{k}=L_{k}X(t_{k})$ with $L_{k}\in{\mathbb{R}}^{m_{k}\times d}$ and $m_{k}\leq d$, $k=1,\dots n$. We will assume the rows of each $L_{k}$ to be linearly independent. Typically reaction rates are unknown and we wish to infer those from the data. Denote the parameter vector containing those rates by $\theta$. Likelihood-based inference for $\theta$ is hampered by the lack of closed-form expressions for the transition probabilities of $X$. However, if the process were observed continuously over time, the problem would be easier. Therefore, it is natural to employ a data-augmentation scheme where we iteratively sample $X$ on $[0,t_{n}]$ conditional on $v_{1},\dots,v_{n}$ and $\theta$ and then update $\theta$ conditional on $X$. In this paper, we primarily focus on the first step, sampling from $\left(X\mid L_{k}X(t_{k})=v_{k},\,k=1,\dots n\right)$. It is a key objective of this paper to show rigorously how this can be done efficiently.
1.3. Approach: conditioning by guiding
Our approach builds on earlier work in Corstanje et al., (2023) for general Markov processes in case of a single observation. Let us highlight the main points. The law of the process $X$, conditioned to be in a given states at fixed times, is obtained through Doob’s $h$-transform. That is, there is a function $h$ that depends on the transition probabilities of $X$ that induces a change of measure such that under the new measure, ${\mathbb{P}}^{h}$, the process is conditioned to hit the observed states at times of observation. Since $h$ is typically unknown, we replace it by a fully tractable function $g$ that itself induces a change of measure to a measure ${\mathbb{P}}^{g}$. Under certain conditions, ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ and
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}}{\,\mathrm{d}{\mathbb{P}}^{g}}(X)=\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{n}}(X).$ (1.1)
Here, $\Psi$ is known in closed form and does not depend on $h$. This result implies that $h(0,x_{0})=g(0,x_{0})\mathbb{E}^{g}\Psi_{t_{n}}(X)$. As in fact $h(0,x_{0})$ is the likelihood, we obtain an unbiased estimator for the likelihood which can for example be used in a pseudo-marginal MCMC method (Andrieu & Roberts, (2009)). A large part of this paper deals with choosing the map $(t,x)\mapsto g(t,x)$ and making the above argument precise. As we show, absolute continuity in (1.1) is not obvious and requires proof.
1.4. Related literature
Statistical inference for chemical reaction processes has received considerable attention over the past decade. In this section we summarise related work, while in the next section we highlight contributions of this paper. Rathinam & Yu, (2021) consider the setting where one observes a subset of the species counts continuously over time and wants to filter the latent species counts. Reeves & Bhat, (2022) parametrise model transition rates by neural networks, while assuming all trajectories are fully continuously observed. Parameter estimation is then done by gradient ascent to maximise the loglikelihood.
In this paper, we consider discrete time observations and therefore the works below are more closely related to our work. Warne et al., (2019) give an introduction to chemical reaction processes and consider estimation for discrete time partially observation with Gaussian noise focussing on Approximate Bayesian Computation. Fearnhead, (2008) and Golightly & Sherlock, (2019) are probably closest related to our approach. The common starting points of these works is a slightly informal computation that reveals how the reaction rate of the chemical reaction processes changes upon conditioning the process on a future observation (we formalise this argument in Equation (3.4)). While the reaction rate for the conditioned process is intractable, it can be approximated and this simply boils down to choosing $g$ as above. Fearnhead, (2008) approximates $g$ using Euler discretisation of the Chemical Langevin Equation (CLE) assuming the process is fully observed without error. Golightly & Sherlock, (2019) consider conditioning on a partial observation corrupted by Gaussian noise. Their choice of $g$ is based on the linear noise approximation to the CLE. This is shown to outperform the approach of Fearnhead, (2008) and earlier work in Golightly & Wilkinson, (2015). Georgoulas et al., (2017) construct an unbiased estimator for the likelihood using random truncations and computation of matrix exponentials. This in turn is used to exploit the pseudo-marginal MCMC algorithm (Andrieu & Roberts, (2009)) for parameter estimation. Building upon this work Sherlock & Golightly, (2023) introduce the minimal extended statespace algorithm and the nearly minimal extended statespace algorithm to alleviate the problem of choosing a proposal distribution for the truncation level, as required in Georgoulas et al., (2017). Alt & Koeppl, (2023) consider the same setting as we do and derive approximations to the filtering and smoothing distributions using expectation propagation.
1.5. Contribution
We extend the result in Corstanje et al., (2023) for a single complete observation to multiple partial observations in the context of chemical reaction processes. Moreover, we discuss a variation of the next reaction algorithm by Gillespie, (1976) for sampling from a class of reaction networks with unbounded time-dependent reaction intensities. The proposed methods fit within the framework of Backward Filtering Forward Guiding (van der Meulen & Schauer, (2020)), drawing strongly on techniques for exponential changes of measure as outlined in Palmowski & Rolski, (2002). This enables to construct a guided process that at any time takes into account all future conditionings.
Compared to Fearnhead, (2008) and Golightly & Sherlock, (2019), we consider the setting of multiple future conditionings (rather than one), without imposing extrinsic noise on the observations. Moreover, we derive the conditioned process and likelihood ratio in (1.1) on path space. Sufficient conditions on $g$ to guarantee absolute continuity are given in Theorem 3.7. It turns out that the choices for $g$ in Fearnhead, (2008) and Golightly & Sherlock, (2019) satisfy the assumptions of this theorem. In numerical examples we show that flexibility in choosing $g$ is particularly beneficial in case some of the components of the chemical reaction process have counts that vary monotonically over time.
1.6. Outline
We introduce stochastic chemical reaction processes in Section 2 and discuss examples that we will study. In Section 3, we describe the approach of conditioning chemical reaction processes by guiding through Doob’s $h$-transform and present conditions on $g$ such that ${\mathbb{P}}^{h}\ll{\mathbb{P}}^{g}$. In Section 4 we consider various choices for $g$. In Section 5 we present methods for simulation of conditioned chemical reaction processes together with numerical illustrations. We end with a discussion section. The appendix contains various proofs.
1.7. Frequently used notation
Throughout, we assume that we have an underlying right-continuous filtered probability space $\left(\Omega,{\mathcal{F}},\{{\mathcal{F}}_{t}\}_{t},{\mathbb{P}}\right)$ such that ${\mathcal{F}}=\bigvee_{t}{\mathcal{F}}_{t}$, $\{{\mathcal{F}}_{t}\}_{t}$ is a right-continuous filtration and all processes are adapted to the filtration. The restriction of a probability measure ${\mathbb{P}}$ or expectation ${\mathbb{E}}$ to ${\mathcal{F}}_{t}$ is denoted by ${\mathbb{P}}_{t}$ or ${\mathbb{E}}_{t}$, respectively. For a Markov process $X$, with law ${\mathbb{P}}$, we denote the measure induced by a Doob $h$-transform using a function $g$ by ${\mathbb{P}}^{g}$. Expectations with respect to ${\mathbb{P}}^{g}$ is denoted by ${\mathbb{E}}^{g}$. The generator of $X$ is denoted by ${\mathcal{L}}$ and we refer to ${\mathcal{A}}=\frac{\partial}{\partial t}+{\mathcal{L}}$ as the space-time generator. Under ${\mathbb{P}}^{g}$, these are ${\mathcal{L}}^{g}$ and ${\mathcal{A}}^{g}$. The notation is summarized in the table below.
Intuition Law Expectation Space-time generator
original Markov process ${\mathbb{P}}_{t}$ ${\mathbb{E}}_{t}$ ${\mathcal{A}}$
transformed process induced by $g$ ${\mathbb{P}}^{g}_{t}$ ${\mathbb{E}}_{t}^{g}$ ${\mathcal{A}}^{g}$
Furthermore, for a stochastic process $X$, we use the notation $X^{t}=\{X(s)\colon s\leq t\}$. Given $L\in{\mathbb{R}}^{m\times d}$ with $m\leq d$ and $v\in{\mathbb{R}}^{m}$, we denote the inverse image of $v$ under $L$ by $L^{-1}v=\{x\in{\mathbb{R}}^{d}:Lx=v\}$. For functions $f_{1}(t,x)$ and $f_{2}(t,x)$ of time and space, we say that $f_{1}\propto f_{2}$ if there exists a differentiable function $\kappa$ of time such that for all $t,x$, $f_{1}(t,x)=\kappa(t)f_{2}(t,x)$. Derivatives with respect to a variable representing $t$, say $\partial/\partial t$ are denoted by $\partial_{t}$. We denote by
$A_{n}=\{L_{k}X(t_{k})=v_{k},\,k=1,\dots,n\}.$
a set of conditionings. If a measure $\mu$ is absolutely continuous with respect to $\nu$, we write $\mu\ll\nu$.
2. Chemical reaction processes
We construct a stochastic process to model the dynamics of the chemical reaction network described in Section 1.1 following Chapter 1 of Anderson & Kurtz, (2015). Let $X$ be a Markov process on ${\mathbb{S}}$ such that the $i$’th component of $X(t)$, $X_{i}(t)$, represents the concentration of species $S_{i}$ at time $t$. A reaction $\ell$ is represented through a difference vector $\xi_{\ell}\in{\mathbb{S}}$ and an intensity $\lambda_{\ell}\colon[0,\infty)\times{\mathbb{S}}\to[0,\infty)$. That is, $\lambda_{\ell}$ and $\xi_{\ell}$ are such that
${\mathbb{P}}\left(X({t+\Delta})-X(t)=\xi_{\ell}\mid{\mathcal{F}}_{t}\right)=\lambda_{\ell}(t,X(t))\Delta+o(\Delta),\qquad\Delta\downarrow 0.$ (2.1)
The jump probabilities specified in (2.1) correspond to a jump process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and exponential jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$. The space-time process $(t,X_{t})$ has infinitesimal generator
${\mathcal{A}}f(t,x)=\partial_{t}f(t,x)+{\mathcal{L}}_{t}f(t,x),$
where
${\mathcal{L}}_{t}f(t,x)=\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)\left[f(t,x+\xi_{\ell})-f(t,x)\right],\qquad f\in{\mathcal{D}}({\mathcal{L}}_{t}),\>x\in{\mathbb{S}},.$ (2.2)
where, the domain ${\mathcal{D}}({\mathcal{L}}_{t})$ of ${\mathcal{L}}_{t}$ consists of all bounded functions $f\colon{\mathbb{S}}\to{\mathbb{R}}$. Throughout, we impose the following assumptions on the network.
Assumption 2.1.
$(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$ and $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ are such that
  1. (2.1a)
    $\lambda_{\ell}(t,x)\geq 0$ for all $t\geq 0$, $x\in{\mathbb{S}}$ and $\ell\in{\mathcal{R}}$.
  2. (2.1b)
    For all $\ell\in{\mathcal{R}}$, $\xi_{\ell}\in{\mathbb{Z}}^{d}$ is such $\lambda_{\ell}(t,x)>0$ implies $x+\xi_{\ell}\in{\mathbb{S}}$ for all $t\geq 0$ and $x\in{\mathbb{S}}$.
  3. (2.1c)
    For all $t\geq 0$ and $x\in{\mathbb{S}}$, $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)<\infty$.
  4. (2.1d)
    For $f$ with finite support in ${\mathbb{S}}$, $t\geq 0$ and ${\mathcal{L}}_{t}$ defined in (2.2), $\lim_{\left\lvert x\right\rvert\to\infty}{\mathcal{L}}_{t}f(x)=0$.
In particular, $x\mapsto\lambda_{\ell}(t,x)$ need not be linear. For a stochastic process $X(t)$, let $T_{K}=\inf\{t:|X(t)|\geq K\}$ and $T_{\infty}=\lim_{K\to\infty}T_{K}$.
Theorem 2.2 (Theorem 1.22 of Anderson & Kurtz, (2015)).
Suppose 2.1 is satisfied and let $X$ be a jump process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and exponential jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$ satisfying
$X(t)=x_{0}+\sum_{\ell\in{\mathcal{R}}}\xi_{\ell}Y_{\ell}\left(\int_{0}^{t}\lambda_{\ell}(s,X(s))\,\mathrm{d}s\right),\qquad 0\leq t<T_{\infty},$ (2.3)
where $(Y_{\ell})_{\ell\in{\mathcal{R}}}$ are independent, unit rate Poisson processes. Set $X(t)=\infty$ if $t\geq T_{\infty}$. Then the solution to (2.3) is the unique minimal solution of the martingale problem for (2.2). That is, there exists a filtration $\{{\mathcal{F}}_{t}\}$ such that for each $f$ with finite support
$f(t,X(t))-\int_{0}^{t}{\mathcal{A}}f(s,X(s))\,\mathrm{d}s$
is a $\{{\mathcal{F}}_{t}\}$-martingale.
2.1. Distribution of reaction times
With each reaction $\ell\in{\mathcal{R}}$, we associate a reaction time $\tau_{\ell}$. If $\lambda_{\ell}$ is constant in time, it follows from (2.1) that $\tau_{\ell}\mid X(s)=x\sim\mathrm{Exp}\left(\lambda_{\ell}(x)\right)$. This implies that the time $\tau=\min_{\ell\in{\mathcal{R}}}\tau_{\ell}$ that the first reaction occurs satisfies $\tau\mid X(s)=x\sim\mathrm{Exp}\left(\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(x)\right)$. If the reaction times are inhomogeneous in time and 2.2 applies, the reaction times satisfy, see e.g. Gillespie, (1992),
${\mathbb{P}}\left(\tau_{\ell}>\Delta\mid X(t)=x\right)=\exp\left(-\int_{t}^{t+\Delta}\lambda_{\ell}(s,x)\,\mathrm{d}s\right),\qquad t\geq 0,\Delta>0,x\in{\mathbb{S}}.$ (2.4)
Similarly the time for the first reaction to occur has distribution (2.4) with intensity function $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}$. Moreover, the probability distribution of the first reaction in any subset $R\subseteq{\mathcal{R}}$ also satisfies (2.4) with intensity function $\sum_{\ell\in R}\lambda_{\ell}$.
2.2. Chemical master equation and chemical Langevin equation
Denote the transition probabilities of (2.3) by $p$. That is $p(s,x;t,y)={\mathbb{P}}\left(X(t)=y\mid X(s)=x\right)$. The Kolmogorov forward equation yields the chemical master equation given by
$\partial_{t}p(s,x;t,y)=\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,y-\xi_{\ell})p(s,x;t,y-\xi_{\ell})-\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,y)p(s,x;t,y),$
with initial condition $p(s,x;s,y)=1\{y=x\}$. It is well-known, see e.g. Li, (2020) that solutions to the chemical master equation can be approximated by the transition density of solutions to the Chemical Langevin Equation (CLE), which is the SDE
$\,\mathrm{d}N(t)=b_{\mathrm{CLE}}(t,N(t))\,\mathrm{d}t+\sigma_{\mathrm{CLE}}(t,N(t))\,\mathrm{d}W_{\ell}(t),$ (2.5)
where (assuming the reactions to be numbered $1,\ldots,p$)
$\begin{split}b_{\mathrm{CLE}}(t,x)&=\sum_{\ell=1}^{p}\lambda_{\ell}(t,x)\xi_{\ell}\\ \sigma_{\mathrm{CLE}}(t,x)&=\begin{bmatrix}\xi_{1}&\ldots&\xi_{p}\end{bmatrix}\sqrt{\mbox{diag}(\lambda_{1}(t,x),\ldots,\lambda_{p}(t,x))}\end{split}$ (2.6)
and $W$ is an independent ${\mathbb{R}}^{d}$-valued Brownian motion.
2.3. Examples
Example 2.3 (Pure death process).
Our simplest example models a population of initial size $x_{0}$ in which an individual dies in a time interval $(t,t+\Delta)$ with probability $c\Delta+o(\Delta)$ for some constant $c>0$ and $\Delta$ small. Such a process is modelled as chemical reaction process with just one specie and one reaction, namely $(t,x)\mapsto cx$ with $\xi=-1$.
Example 2.4 (Gene Transcription and Translation (GTT)).
A stochastic model for the process in which information is encoded in DNA and transferred to mRNA is described in Section 2.1.1 of Anderson & Kurtz, (2015). The basic model considers the three species Gene (G), mRNA (M) and Protein (P) in the set ${\mathcal{S}}=\{G,M,P\}$. We consider four reactions.
  1. (1)
    Transcription: $G\to G+M$ with rate $\kappa_{1}>0$.
  2. (2)
    Translation: $M\to M+P$ with rate $\kappa_{2}>0$.
  3. (3)
    Degradation of mRNA: $M\to\emptyset$ with rate $d_{M}>0$.
  4. (4)
    Degradation of protein: $P\to\emptyset$ with rate $d_{P}>0$.
[…]
where $Y_{1},Y_{2},Y_{3},Y_{4}$ are independent unit rate Poisson processes. A realization of this process can be found in Figure 1.
Figure 1. Realization of (2.7) using $\kappa_{1}=200$, $\kappa_{2}=10$, $d_{M}=25$, $d_{P}=1$ and initial position $x_{0}=(1,50,10)$. Note that the gene count is constant in this process. Therefore it was omitted from the figure.
Example 2.5 (Enzyme kinetics).
The standard model for describing enzyme kinetics, see e.g. Bersani et al., (2008), where a substrate binds an enzyme reversibly to form an enzyme-substrate complex, which can in turn deteriorate into an enzyme and a product. We thus model Substrate (S), Enzyme (E), Enzyme-substrate (SE) and Product (P) in the species set ${\mathcal{S}}=\left\{S,E,SE,P\right\}$ and consider of the following reactions.
$S+E\overset{\kappa_{1}}{\underset{\kappa_{2}}{\rightleftharpoons}}SE\overset{\kappa_{3}}{\longrightarrow}P+E$ (2.8)
Equivalently:
  1. (1)
    $S+E\to SE$ with rate $\kappa_{1}$.
  2. (2)
    $SE\to S+E$ with rate $\kappa_{2}$.
  3. (3)
    $SE\to P+E$ with rate $\kappa_{3}$.
Then the reaction rates corresponding to the above listed three reactions are given by
  1. (1)
    $\lambda_{1}(x)=\kappa_{1}x_{1}x_{2}$ and $\xi_{1}=(-1,-1,1,0)$.
  2. (2)
    $\lambda_{2}(x)=\kappa_{2}x_{3}$ and $\xi_{2}=(1,1,-1,0)$.
  3. (3)
    $\lambda_{3}(x)=\kappa_{3}x_{3}$ and $\xi_{3}=(0,1,-1,1)$.
Here $(x_{1},x_{2},x_{3})$ refer to species counts of $(S,E,SE)$. Interesting aspects of this example are firstly that the fourth component ($P$) only appears in reaction ($3$) where $1$ is added and therefore is monotonically increasing and secondly that there are absorbig states such as $x=(0,x_{2},0,x_{4})$ where the process is killed as all reaction rates are zero. A realization of this process can be found in Figure 2.
Figure 2. Realization of the forward process of 2.5 using $\kappa_{1}=5$, $\kappa_{2}=5$, $\kappa_{3}=3$ and initial position $x_{0}=(12,10,10,10)$.
3. Doob $h$-transform and guided Markov processes
We assume $x_{0}$ to be known and that $v=LX(T)$ is observed, where $L\in{\mathbb{R}}^{m\times d}$ is of full row rank and $m\leq d$. In case $m=d$, we assume without loss of generality that $L=I$. The extension to multiple observations will follow in a straightforward way from this case. We aim to sample $X$ on $[0,T]$, conditional on the event $\{LX(T)=v\}$. Our approach crucially depends on results from Palmowski & Rolski, (2002) on exponential change of measure.
For a suitable $f\colon[0,T)\times{\mathbb{S}}\to{\mathbb{R}}$ define
$D_{t}^{f}=\frac{f(t,X(t))}{f(0,x_{0})}\exp\left(-\int_{0}^{t}\frac{{\mathcal{A}}f}{f}(s,X(s))\,\mathrm{d}s\right),\qquad t\in[0,T).$ (3.1)
Definition 3.1.
We write $g\in{\mathcal{G}}$ if $g\in{\mathcal{D}}({\mathcal{A}})$ is a strictly positive function such that $D_{t}^{g},\,t\in[0,T)$ is a martingale adapted to $({\mathcal{F}}_{t})_{t}$.
Let ${\mathbb{P}}_{t}$ denote the law of the process $X$, restricted to ${\mathcal{F}}_{t}$, let $g\in{\mathcal{G}}$ and define the change of measure
$\,\mathrm{d}{\mathbb{P}}_{t}^{g}=D_{t}^{g}\,\mathrm{d}{\mathbb{P}}_{t},\qquad t\in[0,T).$ (3.2)
By formula (1.2) of Palmowski & Rolski, (2002), the process under ${\mathbb{P}}_{t}^{g}$ has generator
${\mathcal{L}}_{t}^{g}f(x)=\frac{1}{g(t,x)}\left[{\mathcal{L}}_{t}\left(fg\right)(t,x)-f(x){\mathcal{L}}_{t}g(t,x)\right]=\sum_{\ell}\lambda_{\ell}^{g}(t,x)\left(f(x+\xi_{\ell})-f(x)\right),$ (3.3)
with
$\lambda_{\ell}^{g}(t,x)=\alpha^{g}_{\ell}(t,x)\lambda_{\ell}(t,x)$ (3.4)
and
$\alpha^{g}_{\ell}(t,x)=\frac{g(t,x+\xi_{\ell})}{g(t,x)},\qquad(\ell,t,x)\in{\mathcal{R}}\times[0,T)\times{\mathbb{S}}.$ (3.5)
Definition 3.2.
The guided process induced by $g$ on $[0,t]$ is defined as the process $X$ under $\mathbb{P}^{g}_{t}$.
Some authors refer to this process as the twisted process (see e.g. Del Moral & Penev, (2017)). Under ${\mathbb{P}}_{t}^{g}$, the process is a chemical reaction process with the same jumps as under ${\mathbb{P}}_{t}$ but with adjusted jump intensities in (3.4). The jump intensities in (3.4) suggest that the process obtained from the change of measure in (3.1) is invariant under multiplying $g$ by a function of time. This is formalized in 3.3.
Lemma 3.3 (Invariance under time scaling).
Suppose $g\in{\mathcal{G}}$ and let $c\colon[0,T]\to{\mathbb{R}}_{+}$ be a differentiable function. Then $D_{t}^{cg}=D_{t}^{g}$.
Proof.
Since $c$ only depends on time
[…]
Upon denoting the transition probabilities of $X$ by $p$, i.e. ${\mathbb{P}}(X_{T}\in A\mid X_{t}=x)=\sum_{y\in A}p(t,x;T,y)$ for $A\subseteq{\mathbb{S}}$, it follows from Example 2.4 of Corstanje et al., (2023) that the conditioned process $\left(X\mid LX(T)=v\right)$ is obtained by taking $g=h$ with
$h(t,x)=\sum_{\zeta\in L^{-1}v}p(t,x;T,\zeta),$ (3.6)
where $L^{-1}v=\{y\in{\mathbb{S}}:Ly=v\}$. Throughout, we assume $h(0,x_{0})>0$. Note that $h$ satisfies Kolmogorov’s backward equation and therefore $h\in{\mathcal{G}}$ and ${\mathcal{A}}h=0$. Hence,
$\,\mathrm{d}{\mathbb{P}}_{t}^{h}=\frac{h(t,X(t))}{h(0,x_{0})}\,\mathrm{d}{\mathbb{P}}_{t},\qquad t<T.$
Intuitively, ${\mathbb{P}}_{t}^{h}$ gives more mass to paths where $h(t,X(t))$ is large. Absolute continuity of ${\mathbb{P}}_{t}^{h}$ with respect to ${\mathbb{P}}_{t}$ won’t hold under the limit $t\uparrow T$ as under ${\mathbb{P}}^{h}_{t}$ the process is conditioned on the event $\{LX_{T}=v\}$. Unfortunately, $h$ is intractable as the transition probabilities $p$ are only known in closed-form in very specific cases. Therefore, consider $g\in{\mathcal{G}}$ that acts as a tractable substitute for $h$. Then, as a consequence, the process $X$ can be simulated under the measure ${\mathbb{P}}_{t}^{g}$ and moreover, for $t<T$,
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}_{t}}{\,\mathrm{d}{\mathbb{P}}^{g}_{t}}(X)=\frac{D_{t}^{h}}{D_{t}^{g}}=\frac{h(t,X(t))}{g(t,X(t))}\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t}^{g}(X),$ (3.7)
where
$\Psi_{t}(X)=\exp\left(\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right).$ (3.8)
To evaluate $\Psi$, we note that a direct computation yields
$\frac{{\mathcal{A}}g}{g}(s,x)=\partial_{s}\log g(s,x)+\sum_{\ell\in{\mathcal{R}}}\left(\lambda_{\ell}^{g}(s,x)-\lambda_{\ell}(s,x)\right).$ (3.9)
Theorem 3.4 (Limiting case of (3.7) as $t\uparrow T$).
Suppose $h$ is as in (3.6) and $g\in{\mathcal{G}}$ is such both $\frac{\partial}{\partial t}\log g$ and $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ are bounded. Then ${\mathbb{P}}_{T}^{h}\ll{\mathbb{P}}_{T}^{g}$ with
$\frac{\,\mathrm{d}{\mathbb{P}}_{T}^{h}}{\,\mathrm{d}{\mathbb{P}}_{T}^{g}}(X)=\frac{g(0,x_{0})}{h(0,x_{0})}\frac{\Psi_{T}(X)}{g(T,X(T-))}\mathbf{1}_{\{LX(T)=v\}},$ (3.10)
where $X(T-)=\lim_{t\uparrow T}X(t)$ $\mathbb{P}^{g}$-almost surely.
Proof.
First note that, as $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ is bounded. $X$ has finitely many jumps in $[0,T]$ and $X(T-)$ and $X(T)$ exist ${\mathbb{P}}^{g}$-almost surely. By (3.7), ${\mathbb{P}}_{t}^{h}\ll{\mathbb{P}}_{t}^{g}$ for $t<T$ and hence for any ${\mathcal{F}}_{s}$-measurable function $f_{s}$ for $s<T$ and $t\in(s,T)$,
${\mathbb{E}}^{h}_{t}f_{s}(X)={\mathbb{E}}_{t}^{g}\left[f_{s}(X)\frac{g(0,x_{0})}{h(0,x_{0})}\frac{h(t,X(t))}{g(t,X(t))}\exp\left(\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right)\right].$ (3.11)
Now note that as $t\uparrow T$, ${h}(t,X(t))={\mathbb{E}}\left[\mathbf{1}\{LX(T)=v\}\mid{\mathcal{F}}_{t}\right]\to{\mathbb{E}}\left[\mathbf{1}\{LX(T)=v\}\mid{\mathcal{F}}_{T}\right]=\mathbf{1}\{LX(T)=v\}$. Hence, taking $\lim_{t\uparrow T}$ on both sides of (3.11) and interchanging limit and expectation finishes the proof. To see that the swap is allowed, note that it follows from (3.9) that $\frac{{\mathcal{A}}g}{g}$ is bounded. Moreover, as $g$ is differentiable in $t$ (as $g\in{\mathcal{G}}$), the maps $t\mapsto g(t,x)$ are continuous for all $x\in{\mathbb{S}}$ and thus $t\mapsto g(t,X(t))$ is ${\mathbb{P}}^{g}$-almost surely cadlag on $[0,T]$ with finitely many jumps and we can find a positive lower bound for it. ∎
Corollary 3.5.
The likelihood of $x_{0}$ based on the observation $v$ is given by
$h(0,x_{0})=g(0,x_{0}){\mathbb{E}}_{T}^{g}\left[\frac{\Psi_{T}(X)}{g(T,X(T-)}\mathbf{1}_{\{LX(T)=v\}}\right].$ (3.12)
Proof.
This follows immediately upon integrating (3.10) with respect to ${\mathbb{P}}^{g}$. ∎
We present one class of functions that satisfies the conditions of 3.4
Proposition 3.6.
Suppose $g\in{\mathcal{D}}({\mathcal{A}})$ is such that for all $x\in{\mathbb{S}}$ the map $t\mapsto g(t,x)$ is continuously differentiable on $[0,T]$ and there exist $\delta,K$ such that $0<\delta\leq g(t,x)\leq K<\infty$ for all $(t,x)\in[0,T]\times{\mathbb{S}}$. Then, by 3.4, ${\mathbb{P}}_{T}^{h}\ll{\mathbb{P}}_{T}^{g}$.
Proof.
As $g$ is bounded away from zero, Lemma 3.2 of Chapter 4 in Ethier & Kurtz, (2009) implies that $D_{t}^{g}$ is a martingale. Hence $g\in{\mathcal{G}}$. Moreover, $\partial_{t}\log g=g^{-1}\partial_{t}g$ is bounded as $g\geq\delta$ and $\partial_{t}g(t,x)$ is continuous on $[0,T]$ for fixed $x$. Lastly, for all $(t,x)$ we have that $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(t,x)\leq(K/\varepsilon)\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)$, which is bounded by Assumption (2.1c). ∎
The following proposition is useful to numerically evaluate the likelihood of a sampled guided process.
Proposition 3.7.
Suppose $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, $x_{0},\dots,x_{N}\in{\mathbb{S}}$ and
$x(t)=\sum_{j=0}^{N-1}x_{j}\mathbf{1}_{[t_{j},t_{j+1})}(t).$
Then, with $\alpha_{j}(s):=g(s,x_{j+1})/g(s,x_{j})$
$\begin{split}&\frac{g(0,x_{0})}{g(T,x(T-))}\exp\left(\int_{0}^{T}\frac{{\mathcal{A}}g}{g}(s,x(s))\,\mathrm{d}s\right)\\ &\qquad=\left(\prod_{j=0}^{N-2}\frac{1}{\alpha_{j}(t_{j+1})}\right)\exp\left(\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x_{j})\left[\alpha_{j}(s)-1\right]\,\mathrm{d}s\right)\end{split}$
Proof.
We have
$\exp\left(\int_{0}^{T}\partial_{s}\log g(s,x(s))\,\mathrm{d}s\right)$ $=\exp\left(\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\frac{\,\mathrm{d}}{\,\mathrm{d}s}\log g(s,x_{j})\,\mathrm{d}s\right)$
$=\frac{g(t_{N},x_{N-1})}{g(t_{0},x_{0})}\prod_{j=0}^{N-2}\frac{g(t_{j+1},x_{j})}{g(t_{j+1},x_{j+1})}.$
Moreover,
$\int_{0}^{T}\frac{{\mathcal{L}}g}{g}(s,x(s))\,\mathrm{d}s=\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x_{j})\left[\frac{g(s,x_{j+1})}{g(s,x_{j})}-1\right]\,\mathrm{d}s$
The result follows from (3.9) and upon observing that $g(T,x(T-))=g(t_{N},x_{N-1})$. ∎
4. Choices for $g$
What remains is a specification of the maps $(t,x)\mapsto g(t,x)$ satisfying the assumptions of 3.4. In this section, we demonstrate that the results in Fearnhead, (2008) and Golightly & Sherlock, (2019) can be considered as specific specific choices of $g$ satisfying these assumptions. We also propose a more easily computable choice of $g$ which also yields absolute continuity. Lastly we discuss a choice for $g$ that, under one additional assumption on the reaction network, yields equivalence of the measures ${\mathbb{P}}_{T}^{h}$ and ${\mathbb{P}}_{T}^{g}$ (rather than absolute continuity). In the following, $C$ denotes a covariance matrix.
4.1. Choices based on the chemical Langevin equation
Consider the CLE given in (2.5). Set $a_{\mathrm{CLE}}={\sigma_{\mathrm{CLE}}}{\sigma_{\mathrm{CLE}}}^{\prime}$. Fearnhead, (2008) proposed to choose $g$ based on the Euler discretization of the CLE. This gives
$g_{\rm F}(t,x)={\mathcal{N}}\left(v;L(x+b_{\mathrm{CLE}}(t,x)(T-t)),La_{\mathrm{CLE}}(t,x)L^{\prime}(T-t)+C\right),$
Since $g$ is obtained as a composition of continuously differentiable maps, we infer absolute continuity $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g}_{T}$ from 3.6.
Golightly & Sherlock, (2019) proposed to infer $g$ by using the Linear Noise Approximation (LNA) to the CLE. Here, we discuss their specific choice called “LNA with restart”.
For notational convenience, write $b=b_{\mathrm{CLE}}$ and $\sigma=\sigma_{\mathrm{CLE}}$. Given $t<T$ and $x\in{\mathbb{S}}$, we denote by $z_{T\mid(t,x)}$ and $V_{T\mid(t,x)}$ the solutions at time $T$ of the system of ordinary differential equations
$\,\mathrm{d}z_{s\mid(t,x)}$ $={b}(s,z_{s\mid(t,x)})\,\mathrm{d}s$
$\,\mathrm{d}V_{s\mid(t,x)}$ $=\left(V_{s\mid(t,x)}\left(J_{b}(s,z_{s\mid(t,x)})\right)^{\prime}+J_{b}\left(s,z_{s\mid(t,x)}\right)V_{s\mid(t,x)}+{a}\left(s,z_{s\mid(t,x)}\right)\right)\,\mathrm{d}s$
where $s\in[t,T]$, subject to the initial conditions are given by $z_{t\mid(t,x)}=x$ and $V_{t\mid(t,x)}=0$. Here, $J_{b}$ denotes the Jacobian matrix of ${b}$ (having component $(i,j)$ given by $\partial b_{i}/\partial z_{j}$. The guiding term then is given by
$g_{\mathrm{LNAR}}(t,x)={\mathcal{N}}\left(v;Lz_{T\mid(t,x)},LV_{T\mid(t,x)}L^{\prime}+C\right).$ (4.1)
Absolute continuity $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g}_{T}$ again follows from 3.6.
4.2. Choosing $g$ using the transition density of a scaled Brownian motion
For fixed $\sigma\in{\mathbb{R}}^{d\times d}$ define the process $\tilde{X}$ by $\,\mathrm{d}\tilde{X}(t)={\sigma}\,\mathrm{d}W(t)$, where $W$ is a standard Brownian motion. Denote ${a}={\sigma}{\sigma}^{\prime}$ and assume that ${a}$ is such that $L{a}L^{\prime}$ is strictly positive definite. We derive $g$ from backward filtering the process $\tilde{X}$ with observation $V\mid\tilde{X}(T)\sim{\mathcal{N}}(L\tilde{X}(T),C)$. Let $q$ denote the density of the ${\mathcal{N}}(0,C)$-distribution. The density of $V$, conditional on $\tilde{X}(t)=x$ is given by $u(t,x):=\int\tilde{p}(t,x;T,y)q(v-Ly)\,\mathrm{d}y$. It follows from the results in Mider et al., (2021) that $u(t,x)\propto g(t,x)$ where
$g(t,x)=\exp\left(-\frac{1}{2}x^{\prime}H(t)x+F(t)^{\prime}x\right).$ (4.2)
Here, for $t\leq T$, $H(t)$ and $F(t)$ satisfy the ordinary differential equations
[…]
These differential equations can be solved in closed-form:
$H(t)=z(t)H(T)\quad\text{and}\quad F(t)=z(t)F(T),$ (4.4)
with $z(t)=\left(I+H(T){a}(T-t)\right)^{-1}$
Note that we have not yet specified $\sigma$. Depending on its choice, $g$ may be radically different from $h$. Nevertheless, it can be used as a change of measure to condition paths of $X$ on the event $\{LX_{T}=v\}$.
Theorem 4.1.
Let $g$ be defined by (4.2) and (4.3). Then ${\mathbb{P}}_{T}^{h}\ll{\mathbb{P}}_{T}^{g}$ with
$\frac{\,\mathrm{d}{\mathbb{P}}_{T}^{h}}{\,\mathrm{d}{\mathbb{P}}_{T}^{g}}=\frac{g(0,x_{0})}{h(0,x_{0})}\exp\left(\int_{0}^{T}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right)\exp\left(-\frac{1}{2}v^{\prime}C^{-1}v\right)\mathbf{1}_{\{LX(T)=v\}}.$ (4.5)
Proof.
By C.1, $g$ in Equation 4.8 is bounded from above and bounded away from 00. Since $g$ is smooth in $t$ and well-defined in $T$, we clearly have that $g\in{\mathcal{D}}({\mathcal{A}})$ and continuous differentiability of the maps $t\mapsto g(t,x)$ for all $x\in{\mathbb{S}}$ on $[0,T]$. The result thus follows from 3.6. ∎
4.1 is only of interest when the guided process has a positive probability of hitting $v$ at time $T$. This is ensured by 4.2 for which the proof is deferred to Appendix A.
Theorem 4.2.
${\mathbb{P}}^{g}(LX(T)=v)>0$
A convenient choice for the matrix $C$ has computational advantages.
Definition 4.3.
Let $\varepsilon>0$ and set $C=\varepsilon L{a}L^{\prime}$. The corresponding $g$ in (4.2) is denoted by $g_{\varepsilon}$.
Lemma 4.4.
The induced measure $\mathbb{P}^{g_{\varepsilon}}_{T}$ is the law of a chemical reaction process with intensities $\lambda_{\ell}^{g_{\varepsilon}}(t,x)=\alpha^{g_{\varepsilon}}_{\ell}(t,x)\lambda_{\ell}(t,x)$, where
$\alpha^{g_{\varepsilon}}_{\ell}(t,x)=\exp\left(-\frac{d(v,L(x+\xi_{\ell}))^{2}-d(v,Lx)^{2}}{2(\varepsilon+T-t)}\right),$ (4.6)
with $d$ the metric given by
$d(x,y)^{2}=(y-x)^{\prime}(LaL^{\prime})^{-1}(y-x).$ (4.7)
Proof.
It follows from theorems 2.4 and 2.5 in Mider et al., (2021) that if we define
$M(t):=\left(C+L{a}L^{\prime}(T-t)\right)^{-1},$
then $H(t)=L^{\prime}M(t)L$ and $F(t)=L^{\prime}M(t)v$ and therefore
$g(t,x)\propto\exp\left(-\frac{1}{2}(v-Lx)^{\prime}M(t)(v-Lx)\right).$ (4.8)
The proportionality is up to a differentiable time-dependent function and hence does not affect the change of measure by 3.3. The result now follows by substituting this expression into the definition of $\alpha^{g}_{\ell}(t,x)$ and rewriting in terms of the metric $d$. ∎
This implies that only one matrix inverse, $(LaL^{\prime})^{-1}$, needs to be calculated. By Equation 4.6, the intensity of the guided process either becomes very large or small depending on whether $L(x+\xi_{\ell})$ is closer or further away from $v$ than $Lx$ with respect to the metric $d$, which has intuitive appeal. Also note that $\varepsilon$ appears only in the denominator at $T+\varepsilon$, which implies that the choice $C=\varepsilon L{a}L^{\prime}$ conditions the process at time $T+\varepsilon$, thereby precluding explosive behaviour in (4.8) as $t\uparrow T$.
Remark 4.5.
Ideally, we would like to have ${\mathbb{P}}^{g}\left(LX(T)=v\right)=1$. For $g=g_{\varepsilon}$ as in 4.4 this is not guaranteed. Essentially, the choice $C=\varepsilon L{a}L^{\prime}$ yields a process where $LX(t)$ hits $v$ at time $T+\varepsilon$. Since the intensities are bounded from below and above, there is a positive probability of at least one reaction occurring in $(T,T+\varepsilon)$ resulting in $LX(T)\neq v$. In Section 4.3 we discuss the case where $\varepsilon=0$, which does yield ${\mathbb{P}}^{g}(LX(T)=v)=1$.
Remark 4.6.
Comparing $g_{\mathrm{F}}$, $g_{\mathrm{LNAR}}$ and $g_{\varepsilon}$, it is conceivable that $g_{\varepsilon}$ will deviate most from $h$. From a computational point of view $g_{\varepsilon}$ is most attractive, as can be seen from (4.6). Moreover, in Section 5 it turns out that simulation of the guided process is simplified in case of $g_{\varepsilon}$.
Remark 4.7.
In dimension $1$, we derive a better intuition for the choice for $a$ from (4.6) and (4.7). A small choice leads to a higher guided intensity, and thus a process arriving around $v$ quickly and likely to stay near $v$, while a large value of $a$ leads to trajectories that remain unaffected by $\alpha_{\ell}^{g_{\varepsilon}}$ until $t$ approaches $T$. Ideally, we mimic trajectories under ${\mathbb{P}}^{h}$ and thus a good choice of $a$ ensures $g(0,x_{0})\Psi_{T}(X)\mathbf{1}\{LX(T)=v\}/g(T,X(T-))$ is large (with $\Psi_{t}(X)$ defined in Equation (3.8)). This can be verified using Monte-Carlo simulation sampling $X$ under ${\mathbb{P}}^{g_{\varepsilon}}$. Alternatively, a simple choice for $a$ is $a_{\mathrm{CLE}}(0,x_{0})$ or, if we have a complete observation, $a=a_{\mathrm{CLE}}(T,v)$.
4.2.1. Extension to multiple observations
We now extend 4.1 to a result for multiple observations. Consider observations $v_{i}=L_{i}X(t_{i})$, $i=1,\dots,n$ where $0=t_{0}<t_{1}<\cdots<t_{n}$ and assume without loss of generality that $L_{i}\in{\mathbb{R}}^{m_{i}\times d}$ are of full column rank with $m_{i}\leq d$ and $L_{i}=I$ when $m_{i}=d$.
Proposition 4.8.
Let $p$ denote the transition probabilities of $X$ and define for $t\in[t_{k-1},t_{k})$ and $x\in{\mathbb{S}}$
$h(t,x)$ $={\mathbb{P}}\left(L_{i}X(t_{i})=v_{i},\,i=k,\dots,n\mid X(t)=x\right)$ (4.9)
$=\sum_{\zeta_{k}\in L_{k}^{-1}v_{k}}\cdots\sum_{\zeta_{n}\in L_{n}^{-1}v_{n}}p(t,x;t_{k},\zeta_{k})\prod_{i=k}^{n-1}p(t_{i},\zeta_{i};t_{i+1},\zeta_{i+1}).$
Then $h\in{\mathcal{G}}$ and the change of measure (3.2) induces $\left(X\mid L_{k}X(t_{k})=v_{k},\,k=1,\dots,n\right)$.
Proof.
This result is obtained upon following Example 2.4 of Corstanje et al., (2023) using the delta-dirac distribution $\mu(\zeta_{k})=\delta(v_{k}-L_{k}\zeta_{k})$. ∎
We deduce the form of $g$ from Mider et al., (2021) in a similar way compared to a single observation. We consider an auxiliary process $\tilde{X}$ that solves the SDE
$\,\mathrm{d}\tilde{X}(t)={\sigma}(t)\,\mathrm{d}W(t),\qquad X(0)=x_{0},\,t\in[0,t_{n}),$ (4.10)
where ${\sigma}(t)=\sum_{k=1}^{n}{\sigma}_{k}\mathbf{1}_{[t_{k-1},t_{k})}(t)$ and ${a}_{k}={\sigma}_{k}{\sigma}_{k}^{\prime}$ are positive definite $d\times d$ matrices for $k=1,\dots,n$.
For each observation $k$, we consider $V_{k}\mid X(t_{k})\sim{\mathcal{N}}(0,C_{k})$ where $C_{k}$ is an $m_{k}\times m_{k}$ covariance matrix. Suppose $q_{k}$ denote the density of the ${\mathcal{N}}(0,C_{k})$-distribution. Then the transition density of $\tilde{X}$ satisfies for $t\in[t_{k-1},t_{k})$,
$\begin{gathered}\int\tilde{p}(t,x;t_{k},\zeta_{k})q_{k}(v_{k}-L_{k}\zeta_{k})\prod_{i=k}^{n-1}\tilde{p}(t_{i},\zeta_{i};t_{i+1},\zeta_{i+1})q_{i+1}(v_{i+1}-L_{i+1}\zeta_{i+1})\,\mathrm{d}\zeta_{k}\cdots\,\mathrm{d}\zeta_{n}\\ \propto\exp\left(-\frac{1}{2}x^{\prime}H(t)x+F(t)^{\prime}x\right)=:g(t,x).\end{gathered}$ (4.11)
Here, the expressions for $H$ and $F$ can be found by backward solving a system of equations given by
$H(t_{n})$ $=L_{n}^{\prime}C_{n}^{-1}L_{n}$
$F(t_{n})$ $=L_{n}^{\prime}C_{n}^{-1}v_{n}$
and for $t\in(t_{k-1},t_{k})$, $k=1,\dots,n$
$\,\mathrm{d}H(t)$ $=H(t)a_{k}H(t)\,\mathrm{d}t,\qquad H(t_{k})=H_{k}:=L_{k}^{\prime}C_{k}^{-1}L_{k}+H(t_{k}+)$ (4.12)
$\,\mathrm{d}F(t)$ $=H(t)a_{k}F(t)\,\mathrm{d}t,\qquad F(t_{k})=F_{k}:=L_{k}^{\prime}C_{k}^{-1}v_{k}+F(t_{k}+)$
Finally, we let $H$ and $F$ be a right continuous modification of the solution to (4.12), i.e. setting $H(t_{k})=H(t_{k}+)$ and $F(t_{k})=F(t_{k}+)$, $k=1,\dots,n-1$. This system can be solved in closed form: ${H}(t)=z_{k}(t)H_{k}$, ${F}(t)=z_{k}(t)F_{k}$, where, for $k=1,\dots,n$, $z_{k}(t)=\left(I+H_{k}a_{k}(t_{k}-t)\right)^{-1}$.
Remark 4.9.
Alternatively, following the computations in Section 2 of Mider et al., (2021), $g$ can be expressed as
$g(t,x)\propto\exp\left(-\frac{1}{2}\left(v(t)-L(t)x\right)^{\prime}M(t)\left(v(t)-L(t)x\right)\right),$ (4.13)
where, the proportionality is up to a time-dependent function and for $t\in[t_{k-1},t_{k})$
$L(t)=\begin{pmatrix}L_{k}\\ \vdots\\ L_{n}\end{pmatrix}\qquad\text{and}\qquad v(t)=\begin{pmatrix}v_{k}\\ \vdots\\ v_{n}\end{pmatrix}$
and $M(t)=M^{\dagger}(t)^{-1}$ with $M^{\dagger}(t)$ a block matrix with entries
$M^{\dagger}(t)=\begin{pmatrix}C_{i}\mathbf{1}\{i=j\}+\sum_{l=k}^{i\wedge j-1}L_{i}a_{l+1}L_{j}^{\prime}(t_{l+1}-t_{l})+L_{i}a_{k}L_{j}^{\prime}(t_{k}-t)\end{pmatrix}_{i,j=k}^{n},$
$t\in[t_{k-1},t_{k})$. While this representation of $g$ is useful in most proofs, it is computationally more demanding as the matrix dimensions of $H(t)$ and $F(t)$ are $d\times d$, while $M(t)$, $L(t)$ and $v(t)$ have dimensions that increase with the amount of observations.
Using C.2, the proof of 4.1 can be repeated for a result like 4.1 for each observation $k$ by taking limits $t\uparrow t_{k}$ to find that ${\mathbb{P}}_{t_{n}}^{h}\ll{\mathbb{P}}_{t_{n}}^{g}$ with
$\frac{\,\mathrm{d}{\mathbb{P}}_{t_{n}}^{h}}{\,\mathrm{d}{\mathbb{P}}_{t_{n}}^{g}}=\frac{g(0,x_{0})}{h(0,x_{0})}\exp\left(\int_{0}^{t_{n}}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s-\frac{1}{2}\sum_{k=1}^{n}v_{k}^{\prime}C_{k}^{-1}v_{k}\right)\mathbf{1}_{A_{n}}.$ (4.14)
The term $\exp\left(-\frac{1}{2}\sum_{k=1}^{n}v_{k}^{\prime}C_{k}^{-1}v_{k}\right)$ results from evaluating each fraction $g(t_{k},x)/g(t_{k}-,x)$ on the set $\left\{L_{k}x=v_{k}\right\}$. We can repeat 4.2 for each observation to obtain ${\mathbb{P}}^{g}\left(A_{n}\right)>0$. However, similar to 4.5, we generally do not have ${\mathbb{P}}^{g}\left(A_{n}\right)=1$.
4.3. Choosing $g$ so that ${\mathbb{P}}^{h}$ and ${\mathbb{P}}^{g}$ are equivalent
By (4.14), the measures ${\mathbb{P}}^{h}$ and ${\mathbb{P}}^{g}$ are equivalent when ${\mathbb{P}}^{g}\left(A_{n}\right)=1$, as all other terms in the Radon-Nikodym derivative are nonzero. In this section we propose a choice for $g$ which yields equivalence of the measures ${\mathbb{P}}^{h}$ and ${\mathbb{P}}^{g}$ under an additional assumption on the network.
By 4.5, the choice $C=\varepsilon LaL^{\prime}$ can be interpreted as imposing a condition of hiting $v$ at time $T+\varepsilon$. Therefore, intuitively, equivalence can be achieved upon artificially setting all elements in the matrices $C_{k}$ equal to 00. In this case, $g(t_{k},x)$ becomes ill-defined whenever $L_{k}x\neq v_{k}$ and boundedness of $g$ is lost, rendering the earlier proofs invalid. To utilize this choice for $g$ and show equivalence, we build on earlier work in Corstanje et al., (2023).
4.3.1. Setting and absolute continuity
The initial conditions for the differential equations (4.12) for the functions $H$ and $F$ appearing in (4.11) are not defined as the matrices $C_{1},\dots,C_{n}$ are not invertible, but we can still obtain an equivalent of (4.13). Observe that for $t\in[t_{k-1},t_{k})$,
$\begin{gathered}\int\tilde{p}(t,x;t_{k},\zeta_{k})\delta_{v_{k}}(L_{k}\zeta_{k})\prod_{i=k}^{n-1}\tilde{p}(t_{i},\zeta_{i};t_{i+1},\zeta_{i+1})\delta_{v_{i+1}}(L_{i+1}\zeta_{i+1})\,\mathrm{d}\zeta_{k}\cdots\,\mathrm{d}\zeta_{n}\\ \propto\exp\left(-\frac{1}{2}(v(t)-L(t)x)^{\prime}M(t)(v(t)-L(t)x)\right)=:g(t,x),\end{gathered}$ (4.15)
where $L$, $v$ and $M$ are defined in 4.9 but with $C_{1},\dots,C_{n}=0$ and we define $g(t_{k},x)=g(t_{k}+,x)$ for $x\in L_{k}^{-1}v_{k}$. We now formalize one particular property of $g$ that is important for the remainder of the session.
Lemma 4.10.
The map $g$, defined in (4.15), is such that $g\in{\mathcal{G}}$ and for all $k=1,\dots,n$ and $x\in{\mathbb{S}}$,
$\lim_{t\uparrow t_{k}}g(t,x)=\begin{cases}g(t_{k},x)&\text{if }L_{k}x=v_{k}\\ 0&\text{otherwise}\end{cases}.$ (4.16)
and $\partial_{t}\log g(t,x)\leq 0$ for all $(t,x)\in(0,t_{n})\times{\mathbb{S}}$.
Proof.
Through a direct computation using the Schur complement, it can be derived that as $t\uparrow t_{k}$,
$(v(t)-L(t)x)^{\prime}M(t)(v(t)-L(t)x)=\frac{d_{k}(v_{k},L_{k}x)^{2}}{(t_{k}-t)}+o(t_{k}-t),$ (4.17)
where $d_{k}$ is the metric on ${\mathbb{R}}^{m_{k}}$ defined as
$d_{k}(x,y)=\sqrt{(y-x)^{\prime}\left(L_{k}a_{k}L_{k}^{\prime}\right)^{-1}(y-x)},\qquad x,y\in{\mathbb{R}}^{m_{k}}.$ (4.18)
which yields (4.16). Since $v$ and $L$ are piecewise constant and $\frac{\,\mathrm{d}M}{\,\mathrm{d}t}=M(t)\begin{pmatrix}L_{i}a_{k}L_{j}^{\prime}\end{pmatrix}_{i,j=k}^{n}M(t)$ is positive semidefinite, $\partial_{t}\log g(t,x)\leq 0$. ∎
Theorem 4.11.
Let $g$ be defined by (4.15). Then ${\mathbb{P}}^{h}\ll{\mathbb{P}}^{g}$ with
$\frac{\,\mathrm{d}{\mathbb{P}}_{t_{n}}^{h}}{\,\mathrm{d}{\mathbb{P}}_{t_{n}}^{g}}=\frac{g(0,x_{0})}{h(0,x_{0})}\exp\left(\int_{0}^{t_{n}}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right)\mathbf{1}_{A[X^{t_{n}}]},$ (4.19)
where
$A[X^{t_{n}}]=\left\{\sup_{0\leq s<t_{n}}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(s,X(s))<\infty\right\}.$
Proof.
The proof is deferred to Appendix B. ∎
To show that ${\mathbb{P}}^{g}_{t_{n}}(A[X^{t_{n}}])>0$, we require 4.12, combined with the observation that the proof of 4.2 can be repeated with this choice for $g$ upon observing that $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ stays finite on the sets $\{L_{k}x=v_{k}\}$
Proposition 4.12.
$A_{n}\subseteq A[X^{t_{n}}]$
Proof.
The proof is deferred to Appendix B
Remark 4.13.
To see that the reverse inclusion generally does not hold, consider a process conditioned to hit $v$ at time $T$. If $LX(t)=u\neq v$ for $t\in(T-\varepsilon,T]$ where $u$ is such that no reactions exist such that $d(v,L(u+\xi_{\ell}))<d(v,Lu)$, it can be shown that $t\mapsto\frac{g(t,u+\xi_{\ell})}{g(t,u)}$ is bounded and therefore such trajectories are included in $A[X^{T}]$.
4.3.2. Equivalence
(4.19) shows that ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ on the set $A[X^{t_{n}}]$. For simulation purposes, equivalence would be preferable. It follows from (4.19) that this is indeed the case if ${\mathbb{P}}^{g}\left(A[X^{t_{n}}]\right)=1$. This can be shown if the network also satisfies a greedy property.
Assumption 4.14.
For $k=1,\dots,n$, define the metric $d_{k}$ on ${\mathbb{R}}^{m_{k}}$ through (4.18) and suppose that $d_{1},\dots,d_{n}$ are such that for all $k$, $t\in[t_{k-1},t_{k})$ $x\in{\mathbb{S}}\setminus L_{k}^{-1}v_{k}$, there is a reaction $\ell\in{\mathcal{R}}$ such that $\lambda_{\ell}(t,x)>0$ and $d_{k}(v_{k},L_{k}(x+\xi_{\ell}))<d_{k}(v_{k},L_{k}x)$.
Intuitively, 4.14 is satisfied when there is always a reaction available that takes the process closer to the first desired conditioning given being in state $x$ at time $t$.
Theorem 4.15.
Suppose 4.14 is satisfied. Then ${\mathbb{P}}^{g}_{t_{n}}(A_{n})=1$.
Proof.
The proof is deferred to Appendix B
Corollary 4.16.
It follows from 4.12 and 4.15 that, under 4.14, ${\mathbb{P}}^{g}_{t_{n}}(A[X^{t_{n}}])=1$.
The choice for $g$ described in this section has two advantages. The first being that the set $A[X^{t_{n}}]$ on which absolute continuity is obtained is larger than $A_{n}$ and the second being that an additional assumption yields equivalence. However, a representation such as (4.11) is not available in this case. When evaluating (4.11), the matrix products and inversions are computed for matrices of size $d\times d$, while (4.13) requires matrix computations where the size increases with the amount of observations. We thus conclude that from a theoretical point of view we may prefer the construction in this section, but for computational purposes (4.11) should be preferred, especially in case $n$ is large.
4.4. Choosing $g$ to guide a process with monotone components
4.1 and 4.11 prove that choosing $g$ using the transition density of a scaled Brownian motion yields a process that is absolutely continuous with respect to the conditioned process. However, depending on the network, this might not always be a desirable choice.
Consider for instance 2.5. Here, the fourth component of the process (P) is monotonically increasing as it only appears in reaction $3$, where $1$ is added, which can also be seen in Figure 2. Let us for simplicity assume a complete observation at time $T$, that is $X(T)=v_{T}$. Clearly, if $X_{4}(t)=v_{T,4}$ for some $t<T$, reaction $3$ cannot occur anymore for the process to satisfy the conditioning. However, $\lambda_{3}^{g}(t,X(t))\neq 0$ for the choices of $g$ discussed so far. This can lead to trajectories that don’t satisfy the conditioning, which in turn means a low acceptance ratio when sampling.
Alternatively, we choose an auxiliary process $\tilde{X}=(\tilde{Z},\tilde{Y})$, where $\tilde{Z}$ is the ${\mathbb{R}}^{3}$-valued process that solves $\,\mathrm{d}\tilde{Z}(t)=\sigma\,\mathrm{d}W_{t}$ and $\tilde{Y}$ is a homogeneous Poisson process with intensity $\tilde{\theta}$. We stick with the notation $x=(z,y)\in{\mathbb{R}}^{4}$ with $z\in{\mathbb{R}}^{3}$ and $y\in{\mathbb{R}}$, and denote $v_{T}=(z_{T},y_{T})$. Similar to earlier computations, we deduce that
$g(t,(z,y))=\exp\left(-\frac{d(z_{T},z)^{2}}{2(T-t)}\right)\frac{\left(\tilde{\theta}(T-t)\right)^{y_{T}-y}}{(y_{T}-y)!}\exp\left(-\tilde{\theta}(T-t)\right),$ (4.20)
where
$d(z_{T},z)=\sqrt{(z_{T}-z)^{\prime}a^{-1}(z_{T}-z)}$
Clearly,
$\lim_{t\uparrow T}g(t,(z,y))=\begin{cases}1&\text{if }(z,y)=(z_{T},y_{T})\\ 0&\text{otherwise }.\end{cases}$
Since also $\partial_{t}\log g\leq 0$, 4.10 is also holds for this choice of $g$, meaning 4.11 still applies. However, now $\lambda_{3}^{g}(t,x)=0$ whenever $x_{4}=v_{T,4}$.
5. Simulation methods
In this section we discuss methods for simulating the guided process on $[0,T]$. This is a Markov jump process with time-dependent intensity which can become large when time is close to $T$. Simulating a Markov jump process is easy when intensities do not depend on time. In this case, given a state $x$ at time $t$, we simply simulate reaction times $\tau_{\ell}\sim\mathrm{Exp}(\lambda_{\ell}(x))$ for each reaction, set ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ and move $t\leftarrow t+\tau_{\hat{\ell}}$ and $X(t+\tau_{\hat{\ell}})\leftarrow X(t)+\xi_{\hat{\ell}}$. When ${\mathcal{R}}$ contains many reactions, one could alternatively use Gillespie’s algorithm, see e.g. Gillespie, (1976, 1977), which only samples the first reaction time and then searches for the most likely reaction at that time.
To extend this method to chemical reaction processes with time-dependent intensities, we have to sample reaction times whose distribution satisfies (2.4), for which there is generally no direct sampling method. We therefore apply a Poisson thinning step, which results in in Algorithm 1.
Input: $x\in{\mathbb{S}}$ and $t\geq 0$ and an upperbound $\bar{\lambda}_{\ell}$ for $\lambda_{\ell}(\cdot,x)$.
Result: The next reaction $\hat{\ell}\in{\mathcal{R}}$ and the corresponding reaction time $\tau_{\hat{\ell}}$ from the state $(t,x)$.
1 for $\ell\in{\mathcal{R}}$ do
2       Set $t^{*}=t$ ;
3       Sample $\tau_{\ell}^{*}\sim\mathrm{Exp}\left(\bar{\lambda}_{\ell}\right)$ and set $t^{*}\leftarrow t^{*}+\tau_{\ell}^{*}$;
4       Sample $U\sim\mathrm{Unif}(0,1)$ ;
5       if $U\leq{\lambda_{\ell}(t^{*},x)}/{\bar{\lambda}_{\ell}}$ then
6             Accept, set $\tau_{\ell}=t^{*}-t$ ;
7
8      else
9             Reject and return to line 1 ;
10
11       end if
12
13 end for
14Set ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ ;
return $\hat{\ell}$ and $\tau_{\hat{\ell}}$.
Algorithm 1 Next reaction method for time-inhomogeneous rates with a Poisson thinning step
5.1. Simulation of the guided process for underlying processes with time-homogeneous intensities
In many applications, such as the examples in 2.3, the intensities $\lambda_{\ell},\,\ell\in{\mathcal{R}}$ of the underlying process only depend on the state $x$ and don’t have a direct dependence on $t$. Here we consider this case and present a method for simulating the guided process in such a scenario. Note that if the underlying rates are time-dependent, but bounded, these upper bounds can easily be included in the thinning step.
Simulation of a guided process $X\sim\mathbb{P}^{g}$ requires sampling of reaction times with intensities $\lambda^{g}_{\ell}$ given in Equation (3.4). Algorithm 1 is applicable in this case using the bounds from C.1. However, these bounds are quite rough and therefore the acceptance probability is very low, making the algorithm very inefficient.
5.1.1. Special case: guided process induced by $g_{\varepsilon}$
In case we take $g_{\varepsilon}$, the map $t\mapsto\alpha_{\ell}^{g_{\varepsilon}}(t,x)$ (defined in Equation 4.6) is monotone. Therefore, given $(t,x)$, a tight upper bound can be found at either $(t,x)$ or $(T,x)$. Unfortunately, the upper bound at $T$ is typically far from sharp hampering efficiency in simulation, especially when $\varepsilon$ is chosen to be small. To resolve this issue, we employ Algorithm 2. In lines 22, $\lambda_{\ell}(\cdot,x)$ is increasing in time. However, we can choose $\delta$ such that $t+\delta<T$ and use that, on the interval $[t,t+\delta]$, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is upper bounded by $\lambda_{\ell}^{g_{\varepsilon}}(t+\delta,x)$. Hence, on $[t,t+\delta]$, we apply the thinning property for an inhomogeneous Poisson process with rate $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ and if no reaction occurs in this interval, we move $t\leftarrow t+\delta$.
Input: $x\in{\mathbb{S}}$ and $t\geq 0$.
Result: The next reaction $\hat{\ell}\in{\mathcal{R}}$ and the corresponding reaction time $\tau_{\hat{\ell}}$ from the state $(t,x)$.
1 Set +=${\mathcal{R}}_{+}=\{\ell\in{\mathcal{R}}\mid\lambda_{\ell}(x)>0\}$ as the set of possible reactions to occur ;
2 for $\ell\in{\mathcal{R}}_{+}$ do
3       if $s\mapsto\frac{g(s,x+\xi_{\ell})}{g(s,x)}$ is decreasing in $s$ on $[t,T)$ then
4             Apply lines 1-1 of Algorithm 1 with $\bar{\lambda}_{\ell}=\lambda_{\ell}^{g}(t,x)$;
5
6      else
7             Set $t^{*}=t$ and $t_{\mathrm{start}}=t$ ;
8             Choose $\delta>0$ such that $t+\delta<T$ ;
9             Sample $\tau_{\ell}^{*}\sim\mathrm{Exp}\left(\lambda_{\ell}^{g}(t+\delta,x)\right)$ and set $t^{*}\leftarrow t^{*}+\tau_{\ell}^{*}$ ;
10             Sample $U\sim\mathrm{Unif}(0,1)$ ;
11             if $t^{*}>t+\delta$ then
12                   Set $t\leftarrow t+\delta$ and return to line 2
13            else
14                   if $U\leq\lambda_{\ell}^{g}(t^{*},x)/\lambda_{\ell}^{g}(t+\delta,x)$ then
15                         Accept and set $\tau_{\ell}=t^{*}-t_{\mathrm{start}}$;
16
17                  else
18                         Reject and return to line 2
19                   end if
20
21             end if
22
23       end if
24
25 end for
26Set ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}_{+}}\tau_{\ell}$ ;
return $\hat{\ell}$ and $\tau_{\hat{\ell}}$.
Algorithm 2 Next reaction method with a Poisson thinning step for guided rates where the original intensity is independent of time
5.1.2. Choice of $\delta$
In Algorithm 2, the choice for $\delta$ in line 2 is arbitrary. In the simulations, we opted to include a minimum acceptance rate. The acceptance rate of a proposed time $\tau_{\ell}$ is given by
$\eta=\frac{\lambda_{\ell}^{g_{\varepsilon}}(t+\tau_{\ell},x)}{\lambda_{\ell}^{g_{\varepsilon}}(t+\delta,x)}$ (5.1)
A direct computation yields that
$\eta=\exp\left(\frac{d(v,L(x+\xi_{\ell}))^{2}-d(v,Lx)^{2}}{2(T+\varepsilon-t-\delta)}-\frac{d(v,L(x+\xi_{\ell}))^{2}-d(v,Lx)^{2}}{2(T+\varepsilon-t-\tau_{\ell})}\right),$ (5.2)
Upon solving (5.2) for $\delta$, we find that
$\delta=T+\varepsilon-t-\left(\frac{2\log\eta}{d(v,L(x+\xi_{\ell}))^{2}-d(v,Lx)^{2}}+\frac{1}{T+\varepsilon-t-\tau_{\ell}}\right)^{-1}$
Now $\tau_{\ell}$ is not known when choosing $\delta$, but since $\tau_{\ell}\geq 0$, we obtain the desired acceptance rate $\eta$ by choosing
$\delta\geq T+\varepsilon-t-\left(\frac{2\log\eta}{d(v,L(x+\xi_{\ell}))^{2}-d(v,Lx)^{2}}+\frac{1}{T+\varepsilon-t}\right)^{-1}.$ (5.3)
5.2. Simulation studies
Julia written code for the simulation examples of this section is available in Corstanje, (2023).
5.2.1. Death model and comparison to Golightly & Sherlock, (2019)
We consider the pure death process of 2.3 and estimate the distribution of $X(T)$. For convenience, we denote by $g^{v}$ a guiding term $g$ chosen for the conditioning $X(T)=v$. By 3.5, for any $v\in{\mathbb{S}}$
$p(0,x_{0};T,v)=g^{v}(0,x_{0}){\mathbb{E}}^{g^{v}}\left[\frac{\Psi_{T}(X)}{g^{v}(T,X(T-))}\mathbf{1}_{\{X(T)=v\}}\right]$
Hence, for large $N$, we can estimate $p(0,x_{0};T,v)$ by sampling $X_{1},\dots,X_{N}$ from ${\mathbb{P}}^{g^{v}}$ using Algorithm 2 and computing
$\hat{p}(v)=\frac{1}{N}\sum_{i=1}^{N}g^{v}(0,x_{0})\frac{\Psi_{T}(X_{i})}{g^{v}(T,X_{i}(T-))}\mathbf{1}_{\{X_{i}(T)=v\}}$ (5.4)
It is a well-known result that $X(T)\sim\mathrm{Binom}\left(x_{0},e^{-cT}\right)$, so we will use this density for comparison.In the following experiment, we use (5.4) to estimate the density of the $\mathrm{Binom}\left(x_{0},e^{-cT}\right)$-distribution with $x_{0}=50$, $T=1$ and $c=1/2$ with the functions $g=g_{\mathrm{LNA}}$ from the LNA method with restart, see (4.1), $g=g_{\varepsilon}$ chosen using a scaled diffusion, see (4.4), and $g$ chosen using the density of a unit rate poisson process with intensity $\theta$. Given, $v$, the value for $a=\sigma^{2}$ (appearing in $g_{\varepsilon}$) is chosen around $2.5(x_{0}-v)$, which was rougly the optimum of the map $a\mapsto g^{v}(0,x_{0})\Psi_{T}(X)\mathbf{1}_{\{X(T)=v\}}/g^{v}(T,X(T-))$ for the $1\%$, $50\%$ and $99\%$-quantiles of $X(T)$ (see Remark 4.7). For $\theta$, we applied Theorem 4.2 of Corstanje et al., (2023) which proves equivalence of the measures when $\theta\leq\min\{cx:x\in\{x_{0},x_{0}-1,\dots,v\}\}=cv$ and chose $\theta=cv$; the largest allowed value. For both the diffusion and the LNA guiding term, we used $C=\varepsilon=10^{-5}$.
For forward simulation under ${\mathbb{P}}^{g}$ we employ Algorithm 2 where, for the diffusion, $\delta$ is chosen according to (5.3). For the LNA and poisson $g$, such a closed form expression cannot be derived, and we thus, given $t<T$, set $\delta=\frac{T-t}{2}$. This does not affect the validity of the algorithm, only the acceptance probability of a sampled reaction time. The estimated probability mass functions as well as the percentages of paths with $X(T)=v$ for each $v$ using (5.4) with $N=15000$ are plotted in Figure 3. The sum of squared errors $\sum_{k}(q(v)-\hat{p}_{v})^{2}$, with $q$ denoting the probability mass function of the $\mathrm{Binom}(x_{0},e^{-cT})$-distribution, for each method is shown in table Table 1.
Figure 3. Left: Estimates of the probability mass function of $X(T)\mid X(0)=x_{0}$ using the guiding functions $f$ from the LNA method, a scaled diffusion and a Poisson process. The upper barplot is the true density. For each $v$, we estimated using (5.4) with $N=15000$. Right: The percentages of sample paths with $X(T)=v$ for each $v$ for the three methods.
Method Squared error
LNA (with restart) $1.6\cdot 10^{-4}$
Diffusion guiding term $4.6\cdot 10^{-4}$
Poisson guiding term $1.1\cdot 10^{-5}$
Table 1. sum of squared errors of the estimates from Figure 3 for each of the methods
From Figure 3 and Table 1, we first observe that the panel on the left hand side of Figure 3 appears to confirm $\mathbb{P}_{T}^{h}\ll\mathbb{P}_{T}^{g}$ for all choices of $g$. Moreover, we see that the overall error of the Poisson guiding term is lowest and the diffusion guiding term performs particularly well for low values of $v$ for which the process has to make a lot of jumps. However when conditioning on high values, the amount of sample paths that end up in the right point is low, leading to a larger variance in the estimator. The performance of the Poisson guiding term is explained due to the typically lower guiding term, which is 00 when the conditioning is reached while the other two guiding terms can go past it. Since the measures are equivalent for the Poisson, all processes end up in $v$.
5.2.2. Gene transcription and translation
We start with the GTT-model presented in 2.4. Figure 4 contains realizations of the process $X$ under ${\mathbb{P}}^{g}$ conditioned to hit $(1,11,56)$ at time $T=1$. We used guiding induced by $g_{\varepsilon}$ with $\varepsilon=10^{-5}$ and $a=a_{\mathrm{CLE}}(0,x_{0})$. The plot contains a random sample of 10 out of 1000 simulated trajectories. All trajectories satisfy $X(T)=(1,11,56)$.
Figure 4. Realization of a guided process starting from $x_{0}=(1,50,10)$ conditioned to be at $(1,10,50)$ at time $T=1$ of the GTT-model from 2.4 with $\kappa_{1}=100$, $\kappa_{2}=10$, $d_{M}=25$ and $d_{P}=1$. The thick line is the original (unconditioned) process.
For Figure 5, we sampled 15 times at random and saved the values of randomly chosen components of a forward simulation to simulate a guided process through multiple partial observations. All of the 1000 trajectories hit all partial observations.
Figure 5. Realization of a guided process through 15 arbitrary observations, with the same settings as in Figure 4.
5.2.3. Enzyme kinetics
As described in 2.5, this is a very interesting example as it contains a monotone component as well as an absorbing state, where the process dies. We use this example to compare various guiding terms. We choose the starting point $x_{0}=(12,10,10,10)$ and parameters $(\kappa_{1},\kappa_{2},\kappa_{3})=(5,5,3)$. We consider three scenarios in which the LNA method with restart, the guiding term obtained from a scaled diffusion presented in Section 4 and the guiding term in which one of the components is replaced by a Poisson guiding term as presented in Section 4.4.
  • Scenario A: The process is conditioned to be at $x_{T,0.01}=(0,15,5,27)$ at time $T=1$. $27$ is the $1\%$ quantile of $X_{4}(T)$, determined through forward simulation.
  • Scenario B: The process is conditioned to be at $x_{T,0.5}=(0,19,1,31)$ at time $T=1$. $31$ is the $50\%$ quantile of $X_{4}(T)$, determined through forward simulation.
  • Scenario C: The process is conditioned to be at $x_{T,0.99}=(0,20,0,32)$ at time $T=1$. $32$ is the $99\%$ quantile of $X_{4}(T)$, determined through forward simulation and $x_{T,0.99}$ is also an absorbing state of the process.
First we simulate 100 trajectories of each process and check the percentage of paths that satisfy $X(T)=x_{T,q}$ for $q=0.01,0.5,0.99$. We used the same $\delta$’s as with the death process. For the diffusion guiding term, we took $a=a_{\mathrm{CLE}}(0,x_{0})$. For the Poisson guiding term, we took for the non-monotone components $3\times 3$ with rows $1-3$ and columns $1-3$ of $a_{\mathrm{CLE}}(0,x_{0})$ for $a$ and $\lambda_{3}(0,x_{0})$; the lower bound of the reaction intensity for the reaction that induces the monotone component for the intensity of the Poisson component. The results are summarized in Figure 6.
Figure 6. Percentages out of $100$ trajectories that satisfy the conditionings set in scenario’s A, B and C.
Just as with the death process in Figure 3, we see that the diffusion guiding term and the LNA guiding term struggle when the monotone component is conditioned not to make many jumps.
6. Discussion
In this paper we have provided sufficient conditions for constructing valid guided processes, where “valid” refers to $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g}_{T}$. We have presented various choices of $g$ and conclude that among those there is no best choices in terms of closeness to true conditioned processes and computational costs. When used within a sequential Monte Carlo or Markov chain Monte Carlo algorithm, a mixture of proposals may be beneficial. The inherent discreteness of chemical reaction processes makes it a hard problem, but it works to our advantage in the sense that guided processes can be constructed on $[0,T]$ where $g(T,x)$ is well defined. This is accomplished by using a guiding term derived from conditioning a diffusion process that is observed with (small) extrinsic noise. The guiding term $g_{\varepsilon}$ has the additional advantage that monotonicity can be exploited for simulating paths ${\mathbb{P}}^{g}$ efficiently.
In the analogous problem of continuous-discrete smoothing for diffusion processes (see e.g. Mider et al., (2021), Beskos et al., (2008), Golightly & Wilkinson, (2008)) there exists a simple random-walk like MCMC-sampler on path space to update guided processes: the preconditioned Crank-Nicolson scheme. Unfortunately, we are not aware of a similar construct for chemical reaction processes or any kernel that leaves the smoothing distribution invariant.
7. Acknowledgments
We thank M. Schauer (Gothenburg University and Chalmers University of Technology, Sweden) for stimulating discussions on the simulation algorithms.
8. Declarations
This work is part of the research project ”Bayes for longitudinal data on manifolds” with project number OCENW.KLEIN.218, which is financed by the Dutch Research Council (NWO). The authors have no competing interests to declare that are relevant to the content of this article.
Appendix A Proof of 4.2
The proof relies on the existence of a chain of reactions that result in the conditioning $LX(T)=v$. We formalize this in A.1.
Definition A.1.
For $x,y\in{\mathbb{S}}$, we say that $y$ can be reached from $x$ within a time interval $(a,b)\subseteq(0,T)$, denoted $x\hookrightarrow y$ in $(a,b)$, if there is a finite collection $\{\ell_{1},\dots,\ell_{I}\}\subseteq{\mathcal{R}}$ of reactions and a partition $U_{1},\dots,U_{I}$ of disjoint intervals for $(a,b)$ such that $y=x+\sum_{i=1}^{I}\xi_{\ell_{i}}$ and for all the set $i$, $\mathrm{supp}\,\lambda_{\ell_{i}}\left(x+\sum_{j<i}\xi_{\ell_{j}}\right)\cap U_{i}$ is of positive Lebesgue measure.
Since we consider a process conditioned on $LX(T)=v$, we can assume there exists an $x_{T}\in{\mathbb{S}}$ such that $Lx_{T}=v$ and $x_{0}\hookrightarrow x_{T}$ in $(0,T)$.
Proof of 4.2.
Notice that, given any $t<T$,
${\mathbb{P}}^{g}\left(X(s)=x_{T}\text{ for all }s\in[t,T]\mid X(t)=x_{T}\right)=\exp\left(-\int_{t}^{T}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(s,x_{T})\,\mathrm{d}s\right).$
Now it can be shown that $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ is bounded through C.1 and Assumption Item  (2.1c) that the right hand side is strictly positive and hence, it suffices to show that ${\mathbb{P}}^{g}\left(\exists t<T:X(t)=x_{T}\right)>0$. In the remainder of this proof, we show this by showing that the probability that the chain of reactions described in A.1 connection $x_{0}$ to $x_{T}$ occurs is positive.
Let $\ell_{1},\dots,\ell_{I}$ be the collection of reactions described in A.1 and let $U_{1},\dots,U_{I}$ be the respective partition of $(0,T)$. For the first reaction to occur, we denote by $T_{\ell}^{(1)}$ the reaction time for each reaction $\ell\in{\mathcal{R}}$ and let $L_{1}$ be the first reaction to occur. That is,
$L_{1}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}T_{\ell}^{(1)},\qquad\text{where}\qquad{\mathbb{P}}^{g}\left(T_{\ell}^{(1)}>t\right)=\exp\left(-\int_{0}^{t}\lambda_{\ell}^{g}(s,x_{0})\,\mathrm{d}s\right).$
Note that,
$\left\{T_{\ell_{1}}^{(1)}\in U_{1},\,T_{\ell}^{(1)}>\sup\,U_{1}\text{ for all }\ell\in{\mathcal{R}}\setminus\{\ell_{1}\}\right\}\subseteq\left\{T_{\ell_{1}}^{(1)}\in U_{1},\,L_{1}=\ell_{1}\right\}$
and thus
$\begin{gathered}{\mathbb{P}}^{g}\left(T_{\ell_{1}}^{(1)}\in U_{1},\,L_{1}=\ell_{1}\right)\geq\\ \left(1-\exp\left(-\int_{U_{1}}\lambda^{g}_{\ell_{1}}(s,x_{0})\,\mathrm{d}s\right)\right)\exp\left(-\int_{U_{1}}\sum_{\ell\in{\mathcal{R}}\setminus\{\ell_{1}\}}\lambda_{\ell}^{g}(s,x_{0})\,\mathrm{d}s\right).\end{gathered}$
Since $\lambda_{\ell_{1}}^{g}(\cdot,x_{0})$ has support in $U_{1}$, the first term is nonzero, while the second term is nonzero by Assumption (2.1c).
For the second reaction, we set
$L_{2}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}T_{\ell}^{(2)}\qquad\text{where}\qquad{\mathbb{P}}^{g}\left(T_{\ell}^{(2)}>t\mid L_{1}\right)=\exp\left(-\int_{T_{L_{1}}^{(1)}}^{t}\lambda_{\ell}^{g}(s,x_{0}+\xi_{L_{1}})\,\mathrm{d}s\right).$
Through similar reasoning, we deduce that
${\mathbb{P}}^{g}\left(L_{2}=\ell_{2},\,T_{\ell_{2}}^{(2)}\in U_{2},\,L_{1}=\ell_{1},\,T_{\ell_{1}}^{(1)}\in U_{1}\right)$
$\quad={\mathbb{P}}^{g}\left(L_{2}=\ell_{2},\,T_{\ell_{2}}^{(2)}\in U_{2}\mid L_{1}=\ell_{1},\,T_{\ell_{1}}^{(1)}\in U_{1}\right){\mathbb{P}}^{g}\left(L_{1}=\ell_{1},\,T_{\ell_{1}}^{(1)}\in U_{1})\right)$
$\quad\geq\left(1-\exp\left(-\int_{U_{2}}\lambda_{\ell_{2}}^{g}\left(s,x_{0}+\xi_{\ell_{1}}\right)\,\mathrm{d}s\right)\right)\exp\left(-\int_{U_{2}}\sum_{\ell\in{\mathcal{R}}\setminus\{\ell_{2}\}}\lambda_{\ell}^{g}(s,x_{0}+\xi_{\ell_{1}})\,\mathrm{d}s\right)$
×$\quad\quad\times\left(1-\exp\left(-\int_{U_{1}}\lambda^{g}_{\ell_{1}}(s,x_{0})\,\mathrm{d}s\right)\right)\exp\left(-\int_{U_{1}}\sum_{\ell\in{\mathcal{R}}\setminus\{\ell_{1}\}}\lambda_{\ell}^{g}(s,x_{0})\,\mathrm{d}s\right).$
Upon iteratively repeating this process, we find
${\mathbb{P}}^{g}\left(L_{i}=\ell_{i},\,T_{\ell_{i}}^{(i)}\in U_{i},\,i=1,\dots,I\right)\geq$
$\prod_{i=1}^{I}\left(1-\exp\left(-\int_{U_{i}}\lambda_{\ell_{i}}^{g}\left(s,x_{0}+\sum_{j<i}\xi_{\ell_{j}}\right)\,\mathrm{d}s\right)\right)$
×$\times\exp\left(-\int_{U_{i}}\sum_{\ell\in{\mathcal{R}}\setminus\{\ell_{i}\}}\lambda_{\ell_{i}}^{g}\left(s,x+\sum_{j<i}\xi_{\ell_{j}}\right)\,\mathrm{d}s\right).$
The right hand size is strictly positive as a finite product of nonzero terms. Therefore ${\mathbb{P}}^{g}\left(\exists t\leq T:X(t)=x_{T}\right)>0$, which finishes the proof. ∎
Appendix B Proofs for Section 4.3
Theorem B.1 (Theorem 3.3 of Corstanje et al., (2023)).
Suppose that there exist a family of ${\mathcal{F}}_{t}$-measurable events $\{A_{j}[X^{t}]\}_{j}$ for each $t\in[0,t_{n})$ so that the following assumptions hold.
  1. (B.1a)
    For all $j$ $s\leq t$, $A_{j}[X^{t}]\subseteq A_{j+1}[X^{t}]$ and $A_{j}[X^{t}]\subseteq A_{j}[X^{s}]$.
  2. (B.1b)
    For all $k$ and $s\in[0,t_{k})$ and $x\in{\mathbb{S}}$,
    $\lim_{t\uparrow t_{k}}{\mathbb{E}}\left[g(t,X(t))\mid X(s)=x\right]={\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\right],$
    Moreover, ${\mathbb{E}}\left[g(t,X(t))\mid X(s)\right]$ is ${\mathbb{P}}$-almost surely bounded for all $t$.
  3. (B.1c)
    For all $k=1,\dots,n$,
    $\lim_{j\to\infty}\lim_{t\uparrow t_{k}}{\mathbb{E}}^{h}_{t}\left[\frac{g(t,X_{t})}{h(t,X_{t})}\mathbf{1}_{A_{j}[X^{t}]}\right]=\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A[X^{t_{k}}]}\right]}{h(0,x_{0})},$
    where $A_{j}[X^{t_{k}}]=\cap_{t\in[0,t_{k})}A_{j}[X^{t}]$, $A[X^{t}]=\bigcup_{j}A_{j}[X^{t}]$ and $A[X^{t_{k}}]=\cap_{t\in[0,t_{k})}A[X^{t}]$.
  4. (B.1d)
    For all fixed $j$, $\Psi_{t}(X)\mathbf{1}_{A_{j}[X^{t}]}$ is ${\mathbb{P}}^{g}$-almost surely uniformly bounded in $t$.
Then for any bounded measurable function $f$,
${\mathbb{E}}\left[f(X)\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\right]}{h(0,x_{0})}\mathbf{1}_{A[X^{t_{k}}]}\right]={\mathbb{E}}^{g}\left[f(X)\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{k}}(X)\mathbf{1}_{A[X^{t_{k}}]}\right]$ (B.1)
for $k=1,\dots,n$. In particular, for $k=n$,
${\mathbb{E}}^{h}\left[f(X)\mathbf{1}_{A[X^{t_{n}}]}\right]={\mathbb{E}}^{g}\left[f(X)\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{n}}(X)\mathbf{1}_{A[X^{t_{n}}]}\right].$ (B.2)
Proof.
The proof of Theorem 3.3 of Corstanje et al., (2023) can be followed with the new limits in (B.1b) and (B.1c) for each of the limits $t\uparrow t_{k}$. ∎
Proof of 4.11.
The proof utilizes B.1 with
$V(t,x)=\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(t,x)\qquad\text{and}\qquad A_{j}[X^{t}]=\left\{\sup_{0\leq s<t}V(s,X(s))\leq j\right\}.$ (B.3)
Then (B.1a) is satisfied by construction. Lemmas B.2-B.4 prove the remaining conditions. ∎
Lemma B.2.
For all $k=1,\dots,n$, $s\in[0,t_{n})$ and $x\in{\mathbb{S}}$,
$\lim_{t\uparrow t_{k}}{\mathbb{E}}\left[g(t,X(t))\mid X(s)=x\right]={\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mid X(s)=x\right].$
Proof.
By dominated convergence and 4.10
$\lim_{t\uparrow t_{k}}{\mathbb{E}}\left[g(t,X(t))\mid X(s)=x\right]$ $=\lim_{t\uparrow t_{k}}\sum_{y\in{\mathbb{S}}}g(t,y)p(s,x;t,y)$
$=\sum_{y\in{\mathbb{S}}}g(t_{k},y)\mathbf{1}\{L_{k}y=v_{k}\}p(s,x;t_{k},y)$
$={\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mid X(s)=x\right].$
Lemma B.3.
For all $k=1,\dots,n$,
$\lim_{j\to\infty}\lim_{t\uparrow t_{k}}{\mathbb{E}}^{h}\left[\frac{g(t,X(t))}{h(t,X(t))}\mathbf{1}_{A_{j}[X^{t}]}\right]=\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A[X^{t_{k}}]}\right]}{h(0,x_{0})}.$
Proof.
Upon defining $Z(t)=g(t,X(t))$, it follows from dominated convergence and 4.10 that for any bounded continuous function $f$,
$\lim_{t\uparrow t_{k}}{\mathbb{E}}f(Z(t))$ $=\sum_{x\in{\mathbb{S}}}\lim_{t\uparrow t_{k}}f\left(g(t,x)\right)p(0,x_{0};t,x)$
$=\sum_{x\in{\mathbb{S}}}f\left(g(t_{k},x)\mathbf{1}\{L_{k}x=v_{k}\}\right)p(0,x_{0};t_{k},x)$
$={\mathbb{E}}f\left(g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\right).$
Hence,
$Z(t)\rightsquigarrow g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\text{ as }t\uparrow t_{k}.$ (B.4)
Now
${\mathbb{E}}\left[Z(t)\mathbf{1}_{A_{j}[X^{t}]}\right]={\mathbb{E}}\left[Z(t)\mathbf{1}_{A_{j}[X^{t_{k}}]}\right]+{\mathbb{E}}\left[Z(t)\mathbf{1}_{A_{j}[X^{t}]\setminus A_{j}[X^{t_{k}}]}\right]$
By (B.4), the first term tends to ${\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A_{j}[X^{t_{k}}]}\right]$ as $t\uparrow t_{k}$. Moreover, $\mathbf{1}_{A_{j}[X^{t}]\setminus A_{j}[X^{t_{k}}]}\downarrow 0$ as $t\uparrow t_{k}$ and thus the second term tends to 00 by Slutsky’s theorem. Hence, by (3.2)
$\lim_{j\to\infty}\lim_{t\uparrow t_{k}}{\mathbb{E}}^{h}\left[\frac{Z(t)}{h(t,X(t))}\mathbf{1}_{A_{j}[X^{t}]}\right]$ $=\frac{1}{h(0,x_{0})}\lim_{j\to\infty}\lim_{t\uparrow t_{k}}{\mathbb{E}}\left[Z(t)\mathbf{1}_{A_{j}[X^{t}]}\right]$
$=\frac{1}{h(0,x_{0})}\lim_{j\to\infty}{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A_{j}[X^{t_{k}}]}\right]$
$=\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A[X^{t_{k}}]}\right]}{h(0,x_{0})}.$
Lemma B.4.
For all $j$,
$\sup_{t<t_{n}}\exp\left(\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right)\mathbf{1}_{A_{j}[X^{t}]}\leq\exp(jt_{n}).$
Proof.
Since ${\mathcal{A}}=\partial_{t}+{\mathcal{L}}$,
$\frac{{\mathcal{A}}g}{g}(s,x)=\partial_{s}\log g(s,x)+\frac{{\mathcal{L}}g}{g}(s,x).$
Note that
$\frac{{\mathcal{L}}g}{g}(s,x)=\sum_{\ell\in{\mathcal{R}}}\frac{\lambda_{\ell}(s,x)\left(g(s,x+\xi_{\ell})-g(s,x)\right)}{g(s,x)}=V(s,x)-\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x)\leq V(s,x).$
By 4.10, $\partial_{t}\log g\leq 0$ and hence, for all $t<t_{n}$ and on $A_{j}[X^{t}]$,
$\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s=\int_{0}^{t}\partial_{s}\log g(s,X(s))\,\mathrm{d}s+\int_{0}^{t}\frac{{\mathcal{L}}g}{g}(s,X(s))\leq jt\leq jt_{n}.$
Proof of 4.12.
By (4.17),
$\lim_{t\uparrow t_{k}}\frac{g(t,x+\xi_{\ell})}{g(t,x)}=\begin{cases}\frac{g(t_{k},x+\xi_{\ell})}{g(t_{k},x)}&\text{if }L_{k}x=L_{k}(x+\xi_{\ell})=v_{k}\\ 0&\text{if }d_{k}(v_{k},L_{k}x)>d_{k}(v_{k},L_{k}(x+\xi_{\ell})\\ \infty&\text{if }d_{k}(v_{k},L_{k}x)<d_{k}(v_{k},L_{k}(x+\xi_{\ell})\end{cases}.$ (B.5)
And therefore
$\lim_{t\uparrow t_{k}}\lambda_{\ell}^{g}(t,x)=\begin{cases}\lambda_{\ell}(t_{k},x)\frac{g(t_{k},x+\xi_{\ell})}{g(t_{k},x)}&\text{if }L_{k}x=L_{k}(x+\xi_{\ell})=v_{k}\\ 0&\text{if }d_{k}(v_{k},L_{k}(x+\xi_{\ell}))>d_{k}(v_{k},L_{k}x)\\ \infty&\text{if }d_{k}(v_{k},L_{k}(x+\xi_{\ell}))<d_{k}(v_{k},L_{k}x)\end{cases}.$ (B.6)
Now, for $t\in[t_{k-1},t_{k})$, define $B_{t}^{k}=\{L_{k}X(s)=v_{k}\text{ for all }s\in[t,t_{k}]\}$. Then it follows that for all $k$ and $t\in[t_{k-1},t_{k})$,
$B_{t}^{k}\subseteq\left\{\sup_{t_{k-1}\leq s<t_{k}}\,\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(s,X(s))<\infty\right\}.$
Hence
$A_{n}=\bigcup_{k=1}^{n}\bigcap_{t_{k-1}\leq t<t_{k}}B_{t}^{k}\\ \subseteq\bigcup_{k=1}^{n}\left\{\sup_{t_{k-1}\leq s<t_{k}}\,\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(s,X(s))<\infty\right\}=A[X^{t_{n}}].$
Proof of 4.15.
Observe that the result follows immediately if we show that for $k=1,\dots,n$, ${\mathbb{P}}^{g}(L_{k}(X(t_{k})=v_{k})>0$. Now note The probability distribution of the next reaction time $\tau$ satisfies, for $t_{k-1}<s<t<t_{k}$ and $x\in{\mathbb{S}}$,
${\mathbb{P}}^{g}\left(\tau>t\mid X(s)=x\right)=\exp\left(-\int_{s}^{t}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(u,x)\,\mathrm{d}u\right).$
Note that in (B.6), the convergence/divergence is at an exponential rate and therefore, by (B.6) and 4.14,
${\mathbb{P}}^{g}\left(\tau>t_{k}\mid X(s)=x\right)=0$
for all $s<t_{k}$ whenever $L_{k}x\neq v_{k}$. Hence, if $L_{k}X(s)\neq v_{k}$ for any $s<t_{k}$, $X$ jumps before time $t_{k}$ with probability $1$. Since this holds for any $s<t_{k}$, $X$ jumps an infinite amount of times if $L_{k}X$ does not hit $v_{k}$. We now consider the jumps that take $L_{k}X$ closer to $v_{k}$ with respect to the metric $d_{k}$. Let
$\tau_{+}=\inf\{t\geq s\colon d_{k}(v_{k},L_{k}X(t))<d(v_{k},L_{k}X(s))\}.$
Then the distribution of $\tau_{+}$ satisfies for any $x$ such that $L_{k}x\neq v_{k}$ and $t_{k-1}<s<t<t_{k}$,
${\mathbb{P}}^{g}(\tau_{+}>t\mid X(s)=x)\leq\exp\left(-\int_{s}^{t}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(u,x)\mathbf{1}_{\{d_{k}(v_{k},L_{k}(x+\xi_{\ell}))<d_{k}(v_{k},L_{k}x)\}}\,\mathrm{d}u\right).$
By 4.14, there are nonzero entries in this summation that tend to $\infty$ exponentially as $u\uparrow t_{k}$ and thus
${\mathbb{P}}^{g}\left(\tau_{+}>t_{k}\mid X(s)=x\right)=0$
for any $s<t_{k}$ and $L_{k}x\neq v_{k}$. In other words, for any $s<t_{k}$ and $L_{k}x\neq v_{k}$,
${\mathbb{P}}^{g}\left(\exists t\in[s,t_{k}]\text{ such that }d_{k}(v_{k},L_{k}X(t))<d(v_{k},L_{k}x)\mid X(s)=x\right)=1.$
Now ${\mathbb{S}}$ is a discrete lattice and thus any set ${\mathbb{S}}\cap K$, where $K\subseteq{\mathbb{R}}^{d}$ is compact, contains only finitely many points. Hence, since $L_{k}X$ keeps jumping with probability $1$ to points strictly closer to $v_{k}$ if it does not hit $v_{k}$, it must hit $v_{k}$ with probability $1$ and thus
${\mathbb{P}}^{g}\left(\exists t\in[s,t_{k}]\text{ such that }L_{k}X(t)=v_{k}\mid X(s)=x\right)=1$
for any $x$ with $L_{k}x\neq v_{k}$. Since this obviously also holds for $L_{k}x=v_{k}$, we have by the law of total probability
${\mathbb{P}}^{g}\left(\exists t\in[s,t_{k}]\text{ such that }L_{k}X(t)=v_{k}\right)=1.$ (B.7)
Now denote the collection of events $B_{s}=\bigcup_{s\leq t\leq t_{k}}\{L_{k}X(t)=v_{k}\}$, $s\in[t_{k-1},t_{k}]$. By (B.7), ${\mathbb{P}}^{g}(B_{s})=1$ for all $s$. Since $B_{s}\downarrow\{L_{k}X(t_{k})=v_{k}\}$ as $s\uparrow t_{k}$, we have by monotonicity of measures that ${\mathbb{P}}^{g}(L_{k}X(t_{k})=v_{k})=\lim_{s\uparrow t_{k}}{\mathbb{P}}^{g}(B_{s})=1$. ∎
Appendix C Additional Lemmas
Lemma C.1.
Let $g$ be defined through (4.8). Then for all $x$, $g(t,x)$ is bounded from below and above uniformly in $t$.
Proof.
Note
$(v-Lx)^{\prime}M(t)(v-Lx)\geq\lambda_{\min}(M(t))\left\lvert v-Lx\right\rvert^{2}=\left(\lambda_{\max}(M^{\dagger}(t))\right)^{-1}\left\lvert v-Lx\right\rvert^{2}.$
Since for all symmetric matrices $A,B$,
$\lambda_{\min}(A)+\lambda_{\min}(B)\leq\lambda_{\min}(A+B)\leq\lambda_{\max}(A+B)\leq\lambda_{\max}(A)+\lambda_{\max}(B),$
we have that
$\left(\lambda_{\max}(M^{\dagger}(t))\right)^{-1}\geq\left(\lambda_{\max}(C)+\lambda_{\max}(LaL^{\prime})(T-t)\right)^{-1}.$
Now $LaL^{\prime}$ is positive definite and thus the right hand side is decreasing in $t$. Hence, a lower bound is given by $\left(\lambda_{\max}(C)+T\lambda_{\max}\left(LaL^{\prime}\right)\right)^{-1}$. We conclude
$g(t,x)\leq\exp\left(-\frac{1}{2}\left\lvert v-Lx\right\rvert^{2}\left(\lambda_{\max}(C)+T\lambda_{\max}\left(LaL^{\prime}\right)\right)^{-1}\right).$
In a similar way, we can show
$g(t,x)\geq\exp\left(-\frac{1}{2}\left\lvert v-Lx\right\rvert^{2}\lambda_{\min}(C)^{-1}\right).$
Corollary C.2 (Multiple observations).
For all $x$, (4.13) is bounded from below and above uniformly in $t$.
Proof.
The proof of C.1 can be repeated using, for $t\in(t_{k-1},t_{k}]$,
$\left\lvert v(t)-L(t)x\right\rvert^{2}=\sum_{j=k}^{n}\left\lvert v_{j}-L_{j}x\right\rvert^{2}\in\left[\left\lvert v_{n}-L_{n}x\right\rvert^{2},\sum_{j=1}^{n}\left\lvert v_{j}-L_{j}x\right\rvert^{2}\right].$